# JEL Classification: C51, C81, D83, O13, Q12, Q54

### v1

@@ 7 Conclusion
There are numerous research questions that economists would like to answer but are unable to because of a lack of data. Recent years have seen a proliferation in the availability of data, in particular remotely sensed earth observation (EO) data. This has allowed economists and other quantitative researchers to be able to answer questions that were previously viewed as unanswerable. However, there is a recency bias in these new data, meaning the past is still a place that remains data poor.
Our study details a new method for overcoming this recency bias in EO data and with this method we conduct an impact evaluation in a data poor environment. To demonstrate this methodology, we use the case of STRV dissemination and adoption in Bangladesh. We combine data from a variety of sources and use innovative approaches to generating ground truth data. Using recent high resolution EO data and deep learning algorithms, we are able to infer where flooding occurred in the past and generate EO data sets at a higher accuracy than previously existed. We combine these new flood and rice area maps, which cover the entire country of Bangladesh for 20 years, with administrative data on seed dissemination efforts in order to conduct a large-scale, long-term evaluation of the effectiveness of STRVs.
Surprisingly, we do not find evidence that STRVs significantly increase EVI values. We confirm this failure to replicate long-standing results from experimental research by using a three year panel of household survey data and more traditional methods for economic impact evaluations. We hypothesize that our failure to replicate is due to the stochastic nature of STRVs, which require flooding of a very specific amount if the seeds are to produce treatment effects greater than zero. We examine this hypothesis by exploring if improvements to our flood metric in terms of duration and intensity. We find that we can produce positive and significant results but only for a narrow and specific set of measures. We buttress this evidence by conducting Monte Carlo simulations using existing experimental data to show just how sensitive results are to mismeasurement.
This paper demonstrates the possibility and challenges in conducting impact evaluations in data poor settings. It also reveals the added dimensionality of the challenge when trying to capture higher order treatment effects associated with a stochastic technology. As the recent economic and geospatial literature shows, and this paper reaffirms, the mismeasurement problem in data can be acute, often biasing or obscure a true signal. However, the methods developed and deployed in this paper open the possibility to answer previously unanswerable economic questions, allowing us to reduce the number of places that remain data poor, including the past.
@@ Appendix A Sub1 Varietal Development and Dissemination
In the paper we briefly summarize the development, agronomic characteristics, and dissemination of STRVs. In this is appendix we expand on those details and provide additional evidence regarding the when and how STRVs will generate non-zero treatment effects relative to non-STRVs.
@@
All STRVs exhibit a similar cellular and molecular mechanism. The Sub1 sequence closely resembles a protein that functions as a transcription factor181818Transcription factors are proteins that regulate genetic expression by binding to specific DNA sequences (Nature-Education, 2014). This transcription factor facilitates the accumulation of the ethylene hormone in response to submergence stress. The accumulation of ethylene also triggers the production of gibberellic acid. The ethylene hormone is critical for the plantβs vegetative growth, while gibberellic acid promotes the elongation of plant shoots. During complete submergence, however, shoot elongation is reduced. Additionally, other plant metabolic processes, such as carbohydrate consumption and chlorophyll breakdown, are slowed down, activating alternative energy pathways. As a result, the plant enters a βhold your breathβ state, conserving energy until the flooding subsides (Emerick and Ronald, 2019). In contrast, normal rice varieties attempt to elongate their stems and leaves to escape deep submergence. This strategy often results in excessive energy expenditure, leaving the plant unable to recover once the submergence is over. So, the flood-escaping strategy for conventional high-yielding rice varieties differs from that of STRVs, making the latter more adaptable and economically viable (IRRI, 2018).
Having isolated Sub1 and understood how it conveys submergence tolerance, researchers at IRRI were then able to use a combination of traditional cross-breeding methods and marker assisted backcross breeding to create Swarna-Sub1, a hybrid with the popular high yielding Swarna rice variety (Yamauchi etΒ al., 2013). Ismail etΒ al. (2013); IRRI (2018); Emerick and Ronald (2019) have estimated that there is no yield penalty for STRVs under normal condition but a yield advantage of around 1 to 3 t/ha during flooding situation. Additionally, Ray etΒ al. (2017) highlight that the Sub1 cultivar out performs after complete submergence at reproductive stage compared to traditional varieties in terms of carbohydrate quantities and total dehydrogenase.
@@
Appendix B Why Not Other Data
@@
In the paper we make the claim that no socioeconomic data exists to allow for a traditional impact evaluation of the long-term, larve scale effects of STRV adoption in Bangladesh. In this appendix we provide evidence to support that claim by discussing the existing nationally or sub-nationally representative surveys, along with their strengths and weakness for answering our research question.
@@
Needs to be panel data
@@
Needs to have observations before and after 2010
@@
Needs to be representative at the national level or at least of rice cultivation
@@
Needs to have data on rice varieties grown and yields
Many candidate data sets satisfied none or only one of the criteria. Frequently a data set would have multiple observations over time but not contain information on rice variety, the critical data need. Only a few satisfied three of the criteria and none satisfied all four. Te best three data sources, other than the RMS data we used in this study, are:
@@
Village Dynamics in South Asia (VDSA): This dataset is an outcome of a research initiative by the International Crop Research Institute for the Semi-Arid Tropics (ICRISAT).191919https://vdsa.icrisat.org/vdsa-database.aspx. Spanning 42 villages in Bangladesh, the dataset comprises survey data from 1,831 households. The survey is designed to be representative of agricultural households in Bangladesh, is a panel, and has data on varieties and yields. However, there is one critical limitation:
@@
Bangladesh Integrated Household Survey (BIHS): This survey was implemented by the International Food Policy Research Institute (IFPRI).202020https://dataverse.harvard.edu/dataverse/IFPRI/?q=BIHS It is nationally representative, a panel, and contains a host of information on plot-level agricultural production and practices, dietary intake of individual household members, anthropometric measurements (height and weight) of all household members, and data to measure womenβs empowerment in agriculture index (WEAI). However, it faces the same limitation as the RMS data set:
@@
The Bangladesh Rice Research Institute-Rice Database (BRRI-RD): As a government-owned rice database it provides comprehensive data across all years and districts regarding total production212121https://bit.ly/47McOjI. It is nationally representative and contains production information. Nevertheless, this dataset is encumbered by certain limitations:
@@
Appendix C Remote Sensing Details
In the paper we provide a non-technical, intuitive summary of the methods used to generate the EO data. In this appendix we provide the technical details sufficient to implement our procedures in alternative contexts or point to the existing publications, data, and code to replicate the deep learning and fushion model work.
C.1 Rice Area Map
In order to train the rice area mapping algorithm, ground truth data first needed to be generated. Three districts were selected to be the base for the study to extrapolate to the country: Barisal, Kurigram, and Rajshahi). The three districts come from the north, central, and south of the country and experience substantial flooding.
@@ C.1 Rice Area Map
In summary, we apply the following process in Google Earth Engine to prepare the data before applying the RF algorithm:
@@
The RF model consists of a 1000 trees, with each tree having a minimum of 5 leaves. We used the Statistical Machine Intelligence and Learning Engine (smile) RF implementation of Google Earth Engine. The random forest is trained with a leave-one-out cross validation scheme, with an additional separated test set. The leave-one-out scheme consists of rotating the districts used for training and leaving one of the districts out each time. The data used here consists of all years except 2020, which is removed and kept as test set. The leave-one-out validation scheme ensures the model is generalizable in space, meaning that it performs similarly outside of the area used for training. The hold out test set of 2020 allows us to validate the model in time, meaning that it will perform similarly outside of the observed years. The accuracies are reported in TableΒ C1.
@@ C.2 Flood Map
As with the rice maps, building reliable flood maps that look so far back into the past raises data availability challenges. While Satellite data allows for analysis geographical areas, particularly in remote or inaccessible regions it suffers from a recency bias, at least in terms of high spatial and temporal resolution products that can see beyond the visible spectrum. For flood mapping, MODIS is often used because of its long history, its availability on a daily basis, and its relatively cloud free image derived from an 8-days composite image. Products like Landsat, which come at a higher spatial resolution and have a longer time series, do not provide such a consistent image series. In particular, Landsat has a low revisit frequency, harmonization issues between Landsat versions, and instrument failure in Landsat 7.
@@
However, MODIS is not without its own limitations. It is an optical sensor that acquires two images daily. But, being an optical sensor, the content of these images is often obscured by clouds. This means that during the tiime of the year when floods are most frequent (the rainy season), MODIS is at its least useful because cloud cover often makes it difficult to see flooding. Furthermore, MODISβs coarse spatial resolution (250m, 500m, and 1km) compared to LANDSAT, makes it unsuitable for accurately extracting flooded information in urban regions or other complex and heterogeneous landscapes. For us, studying the flat rice growing plains in rural Bangladesh, this last issue is not a major concern.
@@
In recent years, radar sensors can overcome the limitations of optical sensors by penetrating cloud cover to consistently detect water signals. Sentinel-1 is an ideal satellite for flood detection as it captures imagines at a high spatial resolution (10m) and is equipped with a radar sensor so that even in cloudy conditions it can capture accurate and consistent flood events. However, Sentinel-1βs weakness, relative to MODIS and Landsat, is that it only became available since 2017 and is thus unsuitable for historical flood mapping.
@@
Given the strengths and weaknesses of each product, we first ran comparisons of food detection between MODIS and Landsat. While Landsat has greater spatial resolution, which is extremely beneficial in pinpointing floods on small rice plots, its coarse temporal resolution was a critical detriment. Landsat often completely missed flood events detected by MODIS because of Landsatβs infrequency of images and the extent of cloud over. Because of this, we decided to use MODIS as are base product for both flood detection and rice mapping.
@@
Having settled on MODIS, we then developed a method to fuse together the shorter time series of Sentinel-1 with the longer MODIS time series as a potential solution to the disadvantages of either sensor. This fusion model allows us to leverage the advantages of high resolution modern flood detection and historical data availability. The data used for building our flood maps are the product of a deep learning model trained on Sentinel-1 derived fractional inundated area to predict historical MODIS satellite data. We utilize a Convolutional Neural - Long Short-Term Memory Network (CNN-LSTM) to take advantage of both the spatial and temporal feature representation of inundation and compare it to a traditional Convolutional Neural Network (CNN). We apply the CNN-LSTM model from 2002 to 2020 over all of Bangladesh at 500m resolution. From this inundation dataset, we extract one map per year with various statistics over the monsoon season: mean, median, max, min, quantiles, day of max and cumulative flooding.
@@
FigureΒ C2 provides tests for accuracy of the fusion model by comparing results with the high resolution Sentinel-1 data. Our model produces results that match the high resolution data with a high degree of accuracy, ranging from .79 to .92. Additional details and validation tests, along with reproducible data and code, has been published by the research team in Giezendanner etΒ al. (2023).
@@
Appendix D Alternative EO Results
In the analysis relying on EO data, we are concerned about two things. First, that our choices regarding key variables (EVI, flooding) are driving the results and that, had we made different choices, our results would not be significant. Second, that our choice regarding the unit of analysis (districts) is driving the results and that, had we conducted the analysis at a more dis-aggregated level, our results would not be significant. To address the first concern, we re-run all of the district-level analysis using three measures of EVI (cumulative, maximum, mean) that differ from the one presented in our main results (median EVI).
As one can see in this appendix, results are not substantially different when we use one of these alternative EVI measures. We also re-run the DID analysis using various cut-offs for flood-prone. The variable of interest in the DID specification was not significant in our main results, and we see this is true for most other definitions of flood-prone. There is, in fact, one definition of flood-prone that produces significant result, but this is one of 42 possible definitions that we test.
To address the second concern, we re-run all of the analysis using upazila (county) as the unit of analysis.222222There are 64 districts in Bangladesh and these are divided into 553 upazilas. We combine many extremely small upazilas in the city and metro area of Dhaka, where rice is not grown, to get a final data set of 503 upazilas. The caveat with using upazilas is that the administrative seed data is at the district-level, so all upazilas in a district share the same year of introduction and same availability of STRV. The event study results are not robust to the choice of the unit of analysis. This may be because while EVI varies from upazila to upazila, the event year is the same for every upazila in the district and we do not consider the degree of flooding within an upazila. When we estimate the DID and TWFE models, upazila-level results reinforce to the district-level results.
Note: All panels show district-level values for the flood measure normalized by the number of years before of after the introduction of STRV seed in that district. Year 0 is the year the seeds first became available in the district. We then plot non-parametric regressions to both the before and after data with confidence intervals.

### v2

@@ 6 Conclusion
There are numerous research questions that economists would like to answer but are unable to because of a lack of data. Recent years have seen a proliferation in the availability of data, in particular remotely sensed earth observation (EO) data. This has allowed economists and other quantitative researchers to answer questions that were previously viewed as unanswerable. However, there is a recency bias in these new data, meaning the past is still a place that remains data poor.
Our study details a new method for overcoming this recency bias in EO data. With this method we conduct an impact evaluation in a data poor environment. To complicate matters, we demonstrate this methodology using a stochastic technology, STRVs, which only have a measurable impact when flooding is βjust right.β We combine data from a variety of sources and use innovative approaches to generating ground truth data. Using recent high resolution EO data and deep learning algorithms, we are able to infer where flooding occurred in the past and generate EO datasets at a higher accuracy than previously existed. We combine these new flood and rice area maps, which cover the entire country of Bangladesh for 20 years, with administrative data on seed dissemination efforts in order to conduct a large-scale, long-term evaluation of the effectiveness of STRVs.
Taking an agnostic approach, we vary the length and duration of the window in which a flood occurs producing a large number of floods that are potentially βjust rightβ for identifying the treatment effects of STRVs. While a majority of these flood metrics are too little or too much flooding to identify significant impacts, we do find a set of flood metrics that produce positive and significant results. These Goldilocks floods produce consistent results across variation in how we measure EVI and our model specification. These results are also robust to excluding regions of the county that face other abiotic stresses and to using household panel data. However, they are not robust to spatially disaggregating the EO data to the smaller upazila-level. We hypothesize that this is because our primary treatment measure - administrative data on STRV seed availability - is measured only at the larger district-level.
This paper demonstrates the possibilities and challenges for conducting impact evaluations in data poor settings. It also reveals the added dimensionality of the challenge when trying to capture higher order treatment effects associated with stochastic technologies. As the recent economic and geospatial literature shows, and this paper reaffirms, the mismeasurement problem in data can be acute, often biasing or obscure a true signal. However, using the methods developed and deployed in this paper we are able to address the Goldilocks Problem and answer previously unanswerable economic questions, allowing us to reduce the number of places that remain data poor, including the past.
@@ Appendix A Sub1 Varietal Development and Dissemination
In the paper, we briefly summarize the development, agronomic characteristics, and dissemination of STRVs. In this Appendix we expand on those details and provide additional evidence regarding the when and how STRVs will generate non-zero treatment effects relative to non-STRVs.
@@
All STRVs exhibit a similar cellular and molecular mechanism. The Sub1 sequence closely resembles a protein that functions as a transcription factor.171717Transcription factors are proteins that regulate genetic expression by binding to specific DNA sequences (Nature-Education, 2014) This transcription factor facilitates the accumulation of the ethylene hormone in response to submergence stress. The accumulation of ethylene also triggers the production of gibberellic acid. The ethylene hormone is critical for the plantβs vegetative growth, while gibberellic acid promotes the elongation of plant shoots. During complete submergence, however, shoot elongation is reduced. Additionally, other plant metabolic processes, such as carbohydrate consumption and chlorophyll breakdown, are slowed down, activating alternative energy pathways. As a result, the plant enters a βhold your breathβ state, conserving energy until the flooding subsides (Emerick and Ronald, 2019). In contrast, normal rice varieties attempt to elongate their stems and leaves to escape deep submergence. This strategy often results in excessive energy expenditure, leaving the plant unable to recover once the submergence is over. So, the flood-escaping strategy for conventional high-yielding rice varieties differs from that of STRVs, making the latter more adaptable and economically viable (IRRI, 2018).
Having isolated Sub1 and understood how it conveys submergence tolerance, researchers at IRRI were then able to use a combination of traditional cross-breeding methods and marker assisted backcross breeding to create Swarna-Sub1, a hybrid with the popular high yielding Swarna rice variety (Yamauchi etΒ al., 2013). Ismail etΒ al. (2013), IRRI (2018), and Emerick and Ronald (2019) have estimated that there is no yield penalty for STRVs under normal condition but a yield advantage of around 1 to 3 t/ha during flooding situation. Additionally, Ray etΒ al. (2017) highlight that the Sub1 cultivar outperforms after complete submergence at reproductive stage compared to traditional varieties in terms of carbohydrate quantities and total dehydrogenase.
@@
Appendix B Pre-Analysis Plan
Using EO data as proxies for on-the-ground truth is common in geography and becoming common in economics. However, for many variables, there is disagreement on the best way to construct EO proxy metrics. This includes the best way to quantify seasonal EVI to proxy for yields and the best way to quantify flooding that is agronomically relevant. This issue is heightened for flooding in our case because of the Goldilocks Problem. In particular, the plausible variation in the flood window creates an opportunity for an unscrupulous researcher to -hack or data mine their way to significant results, testing a large number of flood metrics but only reporting ones that produce the desired outcomes. Our pre-analysis plan OSF was designed to combat the temptation to find and report results from only a subset of proxies that we might try (AlΒ Rafi etΒ al., 2023).181818Note that the PAP was registered on 9 February 2023. While we refer to the third wave of panel data as 2022, this is because we refer to all cropping season data by the year of planting not the year of harvest. Thus data from 2014, 2017, and 2022 all come from the Aman season in which planting occurred in 2014, 2017, or 2022 and harvest at the end of that calendar year, extending into the subsequent calendar year. Data collection for 2022 was not completed until 12 March 2023. Timestamps on the data are available for verification. However, in pre-specifying the analysis, we under-estimated the sensitivity of our results to precisely defined flood measures. All of our pre-specified flood measures produced null results. It was only after the fact that we began to explore this sensitivity and test new flood metrics that were not pre-specified. Results from the original, pre-specified approach are publicly available in our populated pre-analysis plan (Michler etΒ al., 2024).
While pre-analysis plans (PAPs) are common in lab and field experiments, they remain relatively uncommon for observational studies. However, Josephson and Michler (2023) argue that any studying collecting new data, regardless of the type of data, or trying to establish causal effects should develop and register a PAP. They point out that the very first PAP in economics was observational in nature and was created by Neumark (2001), who was studying the impacts of changes to minimum wage on employment - a particularly contentious issue - and one in which a PAP could be developed ahead of an announced wage increase.
Our study, though observational, is well suited to a PAP given that the research relies on a newly collected round of data and that it uses DID and TWFE methods to establish causal effects of STRV adoption. In addition to preventing the -hacking that Brodeur etΒ al. (2020) shows can be common to DID methods, the plan also helps limit research degrees of freedom - the ability of researchers to make many reasonable or justifiable decision during the data cleaning and analysis process. In our context, it is unclear what the best EO proxy is for agronomically relevant flooding or for yield. There are also a large number of reasonable criteria to use for deciding how to deal with outliers, how to convert, scale, and aggregate data, and how to deal with missing values. In the absence of a PAP, a researcher might be tempted to try many alternative options in these domains and then only present the choices that result it strongly significant results. Our PAP ties our hands so that we can avoid the temptation to selective report results favorable to our hypothesis.
The initial draft of this paper, which is available on arXiv, followed the PAP with two substantial deviations. First, we initially planned to measure flooding as the total number of distinct flood events. Since we filed our PAP, the research team developed the fusion model described in Giezendanner etΒ al. (2023) that provides a fractional index of inundation in a pixel. This new approach provides much greater opportunity to develop more precise measures of flooding than the original count of flood events. Second, the event study regression was not pre-specified but was added at the suggestion of Kyle Emerick during a seminar. We view the version of the manuscript on arXiv as our populated pre-analysis plan that reports on all pre-specified results
During the review process, we received recommendations from two anonymous reviewers and the handling editor, Ariel BenYishay, regarding substantial changes to both data and method. These suggestions included dropping the pre-specified IV analysis and focusing results on the large set of potential Goldilocks floods. This involved a major revision and refocus of the paper, with a number of pre-specified results being dropped or relegated to this online appendix. These changes were made in the spirit of Duflo etΒ al. (2020) and Janzen and Michler (2021), who advocate authors to not stick to an inferior pre-specified method when a new, superior method becomes available. In moving away from our pre-specified approach, we have endeavored to be as transparent as possible, which is why we report all regressions that we run in the specification charts in the paper and why we have included an explicit discussion of our dropping our pre-specified analysis to focus on what many would consider exploratory analysis.
Appendix C Why No Other Data
@@
In the paper we make the claim that no socioeconomic data exists to allow for a traditional impact evaluation of the long-term, large scale effects of STRV adoption in Bangladesh. In this Appendix, we provide evidence to support that claim by discussing the existing nationally or sub-nationally representative surveys, along with their strengths and weakness for answering our research question.
@@
must be panel data,
@@
must have observations before and after 2010m
@@
must be representative at the national level or at least of rice cultivation, and
@@
must have data on rice varieties grown and yields.
Many candidate datasets satisfied none or only one of the criteria. Frequently a dataset would have multiple observations over time but not contain information on rice variety, the critical data need. Only a few satisfied three of the criteria and none satisfied all four. The best three data sources, other than the RMS data we used in this study, are:
@@
Village Dynamics in South Asia (VDSA): This dataset is an outcome of a research initiative by the International Crop Research Institute for the Semi-Arid Tropics (ICRISAT).. Spanning 42 villages in Bangladesh, the dataset comprises survey data from 1,831 households. The survey is designed to be representative of agricultural households in Bangladesh, is a panel, and has data on varieties and yields. However, there is one critical limitation:
@@
Bangladesh Integrated Household Survey (BIHS): This survey was implemented by the International Food Policy Research Institute (IFPRI). It is nationally representative, a panel, and contains a host of information on plot-level agricultural production and practices, dietary intake of individual household members, anthropometric measurements (height and weight) of all household members, and data to measure womenβs empowerment in agriculture index (WEAI). However, it faces the same limitation as the RMS dataset:
@@
The Bangladesh Rice Research Institute-Rice Database (BRRI-RD): As a government-owned rice database it provides comprehensive data across all years and districts regarding total production. It is nationally representative and contains production information. Nevertheless, this dataset is encumbered by certain limitations:
@@
C.1 The Rice Monitoring Survey (RMS)
The dataset that closest matches what we require is the Rice Monitoring Survey (RMS), which we use as a robustness check. The RMS is a Gates Foundation-funded project designed to capture varietal turnovers over time in Bangladesh, India, and Nepal. The data was originally collected as two waves of a panel in 2014 and 2017 (Yamano, 2017). Households were selected following a clustered random sampling procedure to ensure the overall survey was representative of rice growing regions of each country. The plan was to collect a third wave in 2020, but data collection efforts were delayed due to COVID until 2022 and only Bangladesh was revisited. In total, 1,500 households were part of the initial sampling frame in Bangladesh. The RMS was able to follow-up with 1,490 households in 2017 and 1,484 households in 2022. This gives us an attrition rate of over a span of eight years.
As the RMS was designed to monitor varietal turnover, the survey instrument is focused on collecting information on rice production at the crop and plot level. While the data is a panel of households, limited information was collected about household well-being. These include demographics, asset ownership, membership with different farming organizations, farmer opinion on the prevalence of different stresses, such as flood, drought, and salinity, and the availability of different rice varieties. No information was collected on consumption, income, or other measures of welfare, such as food security or womenβs empowerment.
We construct an unbalanced panel of households that cultivated rice in Aman season. This gives us a total of 3,865 observations from 1,488 households. Our primary outcome of interest is rice yield at the household level, which we construct by summing up rice harvest (kg) on all plots and dividing by the sum of area (ha) for all rice plot. We consider a household as having adopted STRVs if they report planting any flood tolerant rice variety on at least one plot. To measure the incidence of flooding faced by the household, we combine household GPS location with the EO flood maps we generate.191919While we would ideally use plot locations to match with flood data, only household GPS locations were collected in 2014 and 2017. Thus, we only know the precise location of plots in 2022. We then calculate five measures of flooding for each survey year: cumulative, maximum, mean, median, and a measure of the flood experienced by the village. Our village measure includes cumulative flooding for all households in the village, minus the household in question. We also construct flood time series for each household going back to 2002. In some specifications we use the householdβs historic experience with flooding to calculate the probability of a household experiencing a flood in a given year as an instrument for STRV adoption.
Appendix D Remote Sensing Details
In the paper we provide a non-technical, intuitive summary of the methods used to generate the EO data. In this appendix we provide the technical details sufficient to implement our procedures in alternative contexts or point to the existing publications, data, and code to replicate the deep learning and fusion model work.
D.1 Rice Area Map
To train the rice area mapping algorithm, ground truth data first needed to be generated. Three districts were selected to be the base for the study to extrapolate to the country: Barisal, Kurigram, and Rajshahi). The three districts come from the north, central, and south of the country and experience substantial flooding.
@@ D.1 Rice Area Map
We apply the following process in Google Earth Engine to prepare the data before applying the RF algorithm:
@@
The RF model consists of a 1,000 trees, with each tree having a minimum of 5 leaves. We used the Statistical Machine Intelligence and Learning Engine (SMILE) RF implementation of Google Earth Engine. The random forest is trained with a leave-one-out cross validation scheme, with an additional separated test set. The leave-one-out scheme consists of rotating the districts used for training and leaving one of the districts out each time. The data used here consists of all years except 2020, which is removed and kept as test set. The leave-one-out validation scheme ensures the model is generalizable in space, meaning that it performs similarly outside of the area used for training. The hold out test set of 2020 allows us to validate the model in time, meaning that it will perform similarly outside of the observed years. The accuracies are reported in TableΒ D1.
@@ D.2 Flood Map
As with the rice maps, building reliable flood maps that look so far back into the past raises data availability challenges. While EO data allows for analysis geographical areas, particularly in remote or inaccessible regions it suffers from a recency bias, at least in terms of high spatial and temporal resolution products that can see beyond the visible spectrum. For flood mapping, MODIS is often used because of its long history, its availability on a daily basis, and its relatively cloud free image derived from an 8-days composite image. Products like Landsat, which come at a higher spatial resolution and have a longer time series, do not provide such a consistent image series. In particular, Landsat has a low revisit frequency, harmonization issues between Landsat versions, and instrument failure in Landsat 7.
@@
Further, MODIS is not without its own limitations. It is an optical sensor that acquires two images daily. But, being an optical sensor, the content of these images is often obscured by clouds. This means that during the tiime of the year when floods are most frequent (the rainy season), MODIS is at its least useful because cloud cover often makes it difficult to see flooding. Furthermore, MODISβs coarse spatial resolution (250m, 500m, and 1km) compared to LANDSAT, makes it unsuitable for accurately extracting flooded information in urban regions or other complex and heterogeneous landscapes. For us, studying the flat rice growing plains in rural Bangladesh, this last issue is not a major concern.
@@
In recent years, radar sensors have overcome the limitations of optical sensors by penetrating cloud cover to consistently detect water signals. Sentinel-1 is an ideal satellite for flood detection as it captures imagines at a high spatial resolution (10m) and is equipped with a radar sensor so that even in cloudy conditions it can capture accurate and consistent flood events. However, Sentinel-1βs weakness, relative to MODIS and Landsat, is that it only became available since 2017 and is thus unsuitable for historical flood mapping.
@@
Given the strengths and weaknesses of each product, we first run comparisons of food detection between MODIS and Landsat. While Landsat has greater spatial resolution, which is extremely beneficial in pinpointing floods on small rice plots, its coarse temporal resolution was a critical detriment. Landsat often completely missed flood events detected by MODIS because of Landsatβs infrequency of images and the extent of cloud over. Because of this, we decided to use MODIS as are base product for both flood detection and rice mapping.
@@
Having settled on MODIS, we then developed a method to fuse together the shorter time series of Sentinel-1 with the longer MODIS time series as a potential solution to the disadvantages of either sensor. This fusion model allows us to leverage the advantages of high resolution modern flood detection and historical data availability. The data used for building our flood maps are the product of a deep learning model trained on Sentinel-1 derived fractional inundated area to predict historical MODIS satellite data. We utilize a Convolutional Neural - Long Short-Term Memory Network (CNN-LSTM) to take advantage of both the spatial and temporal feature representation of inundation and compare it to a traditional Convolutional Neural Network (CNN). We apply the CNN-LSTM model from 2002 to 2020 over all of Bangladesh at 500m resolution. From this inundation dataset, we extract one map per year with various statistics over the monsoon season.
@@
FigureΒ D2 provides tests for accuracy of the fusion model by comparing results with the high resolution Sentinel-1 data. Our model produces results that match the high resolution data with a high degree of accuracy, ranging from .79 to .92. Additional details and validation tests, along with reproducible data and code, has been published by the research team in Giezendanner etΒ al. (2023).
@@
Appendix E Remote Sensing Robustness Checks
E.1 District-Level Results Without Coastal Districts
One concern regarding the main results in our analysis is that we might be picking up the impact of other abiotic stressors, and the adoption of seed designed to mitigate those stresses. Key among these is salinity tolerant rice designed to protect against crop loss due to salt intrusion into the water, either through the water table or through coastal flooding. Work by Chen etΒ al. (2017) and Chen and Mueller (2018) measure the impact of saline tolerant rice in Bangladesh. In their work they find very little impact of flooding on rice yields because farmers adjust to floods by delaying planting date. The concern regarding our results is that in coastal regions the effect that we are capturing is the impact of salinity tolerant rice, biasing our estimates of the impact of flood tolerant rice. Absent coastal districts our results might be more muted or their significance may go away completely.
To evaluate the severity of this concern, we conduct our event study, DID, and TWFE analysis on our district-level data but exclude 19 coastal districts where salinity may be a problem. These 19 include all districts in the division of Barisal plus Chandpur, Chittagoing, Coxβs Bazar, Feni, Lakshmipur, and Noakhali (Chittagong division); Gopalganj and Shariatpur (Dhaka division); and Bagerhat, Jessore, Khulna, Narail, and Sakhira (Khulna division).
FigureΒ E1 presents results from the event study regressions. These are indeed more muted in their measured impact of STRVs. For cumulative, maximum, and mean EVI there appears to be no upward trend except in districts that have had STRVs for years. Only when using median EVI do we see a trend that starts earlier but even here the impact is not significant at standard levels of significance.
FigureΒ E2 presents results from the DID specification. Results using this method are actually stronger when we exclude coastal districts. While in the main results of potential flood metrics produce positive and significant results (using cumulative EVI), here of results are positive and significant. Most importantly for the robustness of our main results, Goldilocks floods come from the same quantiles as before. The majority of positive and significant results are from floods measured using the to quantile. We do however see a few Goldilocks floods at very low quantiles when floods are of very short duration, which we do not see in the main results
Turning to our TWFE regressions, FigureΒ E3 reports results using a binary indicator for if flooding fell within the window and FigureΒ E4 reports results using the number of days in the flood window. Results look remarkably similar to our main results. We even see the same share of positive and significant results using the binary indicator and excluding coastal districts and when we include coastal districts: both . We see fewer positive and significant results using days in the window () then when we include coastal districts ().
Based on the preponderance of evidence, we conclude that our results are driven by the effect of submergence tolerant rice on EVI during floods and not by salinity tolerant rice.
E.2 Upazila-Level Results
A second concern regarding the main results is that our choice regarding the unit of analysis (districts) is driving the results and that, had we conducted the analysis at a more disaggregated level, our results would not be significant. All things being equal, we would prefer to use a smaller, less aggregated unit for analysis. This is both for power reasons (there are over 500 upazilas but only 64 districts) and because averaging over the rice growing area in an entire district likely obscures important variation in EVIβs response to flooding. However, the STRV seed data is only available at the district-level and this drives our choice of districts as the unit of analysis.
FigureΒ E5 presents results from the event study regressions. Relative to the district-level results, the upazila-level results show a stronger impact of STRVs. For all the EVI measures, a strong upward trend begins around six years after the introduction of STRVs. Results are particularly strong for cumulative, mean, and median EVI.
FigureΒ E6 presents results from the DID specification. Results using this method are much weaker at the upazila-level when compared to the district-level. While in the main results of potential flood metrics produce positive and significant results (using cumulative EVI), here less than of results are positive and significant, meaning we cannot discount that those significant results are purely due to chance. The few results that are positive and significant do still come from the same quantiles that produced the Goldilocks floods in our main results. But the clear take away is that our district-level DID results are not robust the disaggregation to the upazila-level.
Turning to our TWFE regressions, FigureΒ E7 reports results using a binary indicator for if flooding fell within the window and FigureΒ E8 reports results using the number of days in the flood window. Results are very similar to the DID results. The vast majority of regressions are not significant or negative and significant. As we mention in the main paper, one reason the non-event study results might not be robust to using upazila-level EO EVI and flood data is that the seed data is still at the district-level. This mismatch of units means that we are trying to explain intra-district variation in EVI using a single district-wide seed value. This type of mismatch can introduce noise into estimates and in fact we do find confidence intervals to be much larger in the upazila-level results. Coefficients of similar size in district-level and upazila-level specification charts have much larger confidence intervals in the upazila-level results than in the district-level results.
Whatever the reason, based on the preponderance of evidence, we conclude that our results are not robust to using district-level seed data but upazila-level EVI and flood data. Our main results should be interpreted with this lack of robustness in mind.